A storage room is 900 cm long. The height of the the length to the width to the height is 8:5:7.
1. What is the difference between the height and width in metres?
2. What is the volume of the storage in cubic metre?

Answer by macston(5194)

A storage room is 900 cm long. The height of the the length to the width to the height is 8:5:7.
1. What is the difference between the height and width in metres?
2. What is the volume of the storage in cubic metre?
H:L:W=8:5:7 ; H=height; L=length=900 cm; W=width
W/L=7/5
W=7L/5
L=900 cm(1 m/100 cm)=9 m
W=7(9m)/5
W=63 m/5
H/L=8/5
H=8L/5
H=8(9m)/5
H=72 m/5
H-W=72m/5-63m/5=9m/5 ANSWER 1: The difference between the height and width in meters is 1.8 meters.
V=L*W*H=9m*12.6m*14.4m=1632.96 cu meters ANSWER 2: The volume of the storage in cubic meters is 1632.96 cubic meters.